OpenStudy (anonymous):

Hey could someone help me with this question please? Simplify each expression. That is, write as a single fraction, fully factored, with no common factors and only positive exponents. a) [f(2+h) - f(2)]/h where f(x) 2/x

OpenStudy (anonymous):

first step is \[f(2)=\frac{2}{2}=1\]

OpenStudy (anonymous):

then \[f(2+h)-f(2)=\frac{2}{2+h}-1\] \[=\frac{2-(2+h)}{2+h}=\frac{-h}{2+h}\]

OpenStudy (anonymous):

ohhh ok thank you so much.

OpenStudy (anonymous):

not done yet

OpenStudy (anonymous):

you have to divide this by h

OpenStudy (anonymous):

the h cancels and you are left with \[\frac{-1}{2+h}\]

OpenStudy (anonymous):

oh that makes sense now thanks a lot, i just joined this site and it helps soo much. Have a good day, and thanks again